Calculus II Spring 2002 MA114:004–006 Russell Brown

• Homework C. Due Wednesday, 30 January 2002. Problem #6 on page 434 of the text. Time will be provided in recitation on Tuesday, 29 January 2002 to work on this problem.

• Quiz #3. Thursday, 31 January 2002. Topic: Inverse trigonometric functions. The quiz will be based on the homework problems from section 6.6.

• Homework §6.8 #7, 9, 21, 23, 25, 27, 29, 91, 93, 95, 99. §7.1 #7, 9, 11, 13, 19, 21, 33, 38, 41, 42, 57. Section 6.8 will be discussed in lecture on Monday, 28 January 2002 and section 7.1 will be discussed on Wednesday, 30 January 2002. Please prepare the problems for §6.8 for recitation on Tuesday, 29 January and §7.1 for Thursday, 31 January 2002.

• The first test will be on Tuesday, 5 February 2002 at 7:30-9:30pm in Chemistry-Physics 139. Please make sure you can solve all of the problems in the homework assignments from sections 6.1-6.6, 6.8 and 7.1. The lettered homework assignments are less relevant to preparing for exams.

• A schedule for tutoring in the Mathskeller, CB065, is available at www.mathskeller.org and then the link for Tutors. A schedule of office hours in Mathskeller may found on the link for Instructors. In addition, you should feel free to use the Mathskeller for studying if you have a free hour between classes.

• The notebook of solutions and handouts is in the math library in the basement of POT.

• Assignments are available on the web from a link at http://www.math.uky.edu/∼rbrown/courses

January 25, 2002